OpenStudy (anonymous):

Find the equation of the line containing the given pair of points. (1/4, -1/2) and (3/4, 2)

OpenStudy (anonymous):

Does anyone know how to set this up

OpenStudy (jhannybean):

equation of a line: \(\rm y-y_1 = m(x-x_1)\) Find your slope first : \((x_1~,~y_1)\rightarrow (1/4~,~ -1/2)\), \((x_2~,~y_2)\rightarrow (3/4~,~2)\) \[m=\frac{y_2-y_1}{x_2-x_1} \therefore ~\text{just plug in your points} \] Use \((x_1~,~y_1) = (1/4~,~-1/2)\) and plug those into your equation too. Tell me what you get. :)

OpenStudy (anonymous):

2.5 for the first points

OpenStudy (jhannybean):

you mean for your slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is this right for the slope

OpenStudy (jhannybean):

Nope.

OpenStudy (anonymous):

I alright I tried it again an got -3/4x

OpenStudy (jhannybean):

What is your \(y_2\) and what is your \(y_1\)?

OpenStudy (jhannybean):

I've stated them in my explanation, all you have to do is match them up.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/4 and -1/2

OpenStudy (jhannybean):

\[(x_1~,~\color{red}{y_1})\rightarrow (1/4~,~ \color{red}{-1/2})\]\[(x_2~,~\color{red}{y_2})\rightarrow (3/4~,~\color{red}2)\]Are you familiar with matching up coordinate points?

OpenStudy (anonymous):

yes it just been a long time since I done them

OpenStudy (jhannybean):

 your first value is always your x-coordinate in a point, the second value is always your y-coordinate point. Always. Remember that.